OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ = 7/8. What is the value of cot θ?

OpenStudy (anonymous):



OpenStudy (anonymous):

theta will be sin inverse of 7/8.... then after that idk..cos we dont do cot

OpenStudy (anonymous):

find the third side via pythagoras it is \[\sqrt{8^2-7^2}=\sqrt{64-49}=\sqrt{15}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

then cotangent is "adjacent over opposite"

OpenStudy (anonymous):

oh so cot is like the opposite of tan

OpenStudy (anonymous):

so its 15/8

OpenStudy (anonymous):

no, it is \(\frac{\sqrt{15}}{7}\)

OpenStudy (anonymous):

the "adjacent " side is \(\sqrt{15}\)and the "opposite" side is \(7\)